OpenStudy (anonymous):

Jorge's rectangular garden was 17 feet long and had a perimeter of 54 feet. He made his garden 5 feet longer. What is the area of his new garden?

OpenStudy (luigi0210):

If I remember correctly, the perimeter of a rectangle is \(P=2l+2w\) and you have 17 ft for length. So you can solve it: \(\large 54=2(17)+2w\) \(54=34+2w\) \(20=2w\) \(10=w\) And the length got extended 5 more feet. 17+5=22 \(\large A=l*w\) \(\large A=22*10\) That should be it :)